OpenStudy (honeymonster):

please help..how do i make x the subject? 2/x+1=y

OpenStudy (anonymous):

just solve for x

OpenStudy (anonymous):

is it \[\frac{2}{x}+1=y \]or\[\frac{2}{x+1}=y\]?

OpenStudy (honeymonster):

the 2nd one..

OpenStudy (anonymous):

then it's 2/(x+1)=y 2=y(x+1) 2/y=x+1 2/y-1=x

OpenStudy (honeymonster):

thank you :)